# Quasinormal Modes and Greybody Factors of Charged Symmergent Black Hole

### v1

@@ Abstract
In this paper, we investigate quasinormal modes (QNMs) and greybody factors within the framework of Symmergent gravity, an emergent gravity model with an $R+R^{2}$ curvature sector. Building on our previous work on static spherically-symmetric solutions [Class.Quant.Grav. 40 (2023) 19, 195003], we explore the effects of the key parameters, including the quadratic curvature term $c_{\rm O}$ and the vacuum energy parameter $\alpha$. For both scalar and electromagnetic perturbations, an increase in $\alpha$ leads to a nearly linear rise in both oscillation frequencies and damping rates. The other parameter $c_{\rm O}$ affects the QNMs spectrum nonlinearly. Additionally, the charge $Q$ of the black hole introduces nonlinear behavior, where higher charges amplify the black hole’s electromagnetic field, resulting in increased oscillation frequencies and faster stabilization. These findings enhance our understanding of charged black hole stability and gravitational wave astrophysics. Further, the analysis of greybody factors reveals that increasing $\alpha$, $c_{\rm O}$, and $Q$ reduces the absorption of radiation, with electromagnetic perturbations reaching maximum absorption at slightly lower frequencies compared to scalar perturbations.
@@ I Introduction
The Standard Model (SM) describes the strong, weak, and electromagnetic interactions as a renormalizable quantum field theory (QFT) wein0 , but does not account for gravity. It is unique to the flat Minkowski spacetime because of difficulties in quantizing the curved metric quant-gravity and in transferring QFTs into curved spacetime weinx2 ; weinx3 . In contrast to full QFTs like the SM, flat spacetime effective QFTs are expected to exhibit a closer connection to curved spacetime. This is because they are nearly classical field theories obtained by integrating out high-frequency quantum fluctuations wein0 ; wein1 , in the framework of both Wilsonian and one-particle-irreducible effective actions polchinski ; burgess .
@@
The SM is endowed with a physical UV cutoff (rather than a regulator) when viewed as an effective QFT wein1 . This UV cutoff denoted by $\Lambda_{\wp}$, represents the scale at which a UV completion becomes relevant. As revealed by the null LHC searches lhc , $\Lambda_{\wp}$ could be at any scale above a TeV. In general, in a Lorentz conserving regime, the loop momentum $\ell^{\mu}$ lies in the interval $-\Lambda_{\wp}^{2}\leq\eta_{\mu\nu}\ell^{\mu}\ell^{\nu}\leq\Lambda_{\wp}^{2}$, and it explicitly breaks the Poincaré (translation) symmetry. In the presence of the UV cutoff $\Lambda_{\wp}^{2}$, scalar mass-squareds obtain corrections proportional to $\Lambda_{\wp}^{2}$, and $\Lambda_{\wp}^{2}$ and $\Lambda_{\wp}^{4}$ terms appear in the loop-corrected vacuum energy wein1 ; veltman . Beyond these UV sensitivities wein1 , gauge bosons also acquire mass-squared terms proportional to $\Lambda_{\wp}^{2}$, and which leads to a breaking of gauge symmetries gauge-break1 ; gauge-break2 . In order to alleviate these unnatural UV sensitivities, the most sensible line of action would be to restore the gauge symmetries by means of the Higgs mechanism Higgs1 ; Higgs2 ; Higgs3 . A hindrance to this proposal is the stark difference between the intermediate vector boson mass (Poincaré-conserving) wein2 and the loop-induced gauge boson mass (Poincaré-breaking) dad2023 . Indeed, in the former, the vector boson mass is promoted to a scalar field, which leads to the usual Higgs mechanism Higgs1 ; Higgs2 ; Higgs3 . In the latter, however, it is necessary to first find a Poincaré-breaking Higgs field. In this regard, it turns out that the affine curvature, which is independent of the metric tensor and its connection, is the sought-after Higgs field in that it promotes the UV cutoff $\Lambda_{\wp}$ affine1 ; affine2 and condenses to the usual metrical curvature in the minimum of the metric-affine action affine3 ; affine4 . The defusion of the unnatural UV sensitivity, the restoration of gauge symmetries, the emergence of gravity, and the appearance of new particles without the necessity to couple directly to SM particles are the salient outcomes of this condensation. This whole mechanism is referred to as gauge symmetry-restoring emergent gravity or symmergent gravity in brief. Symmergent gravity has been constructed and applied through various stages dad2023 ; dad2021 ; dad2019 ; dad2024 , starting with the nascent idea of gauge symmetry restoration by means of curvature dad2016 . Some applications of Symmergent gravity can be found in irfanaliben ; Symmergent-bh2 ; Pulice:2024wjg ; Pulice:2023dqw ; dark-sector ; Symmergent-bh2 ; reggiealiben ; cqgirfan ; bh4 ; inf1 .
@@
Black holes are not isolated cosmic bodies; rather, they are dynamic entities that actively interact with their surroundings, profoundly influencing the fabric of spacetime. When a black hole is perturbed, spacetime itself ripples, creating waves that eventually die down as the system settles back to equilibrium. Understanding how these systems respond to small disturbances is a fundamental aspect of physics, providing deep insights into the nature of spacetime and gravity. The study of black hole perturbations began with pioneering work by Regge:1957td , and was significantly advanced by others Vishveshwara:1970cc ; Zerilli:1970se ; Zerilli:1970wzz ; Zerilli:1974ai ; Moncrief:1975sb ; Teukolsky:1972my , leading to Chandrasekhar’s seminal contributions Chandrasekhar:1985kt ; Chandrasekharbook . At the heart of this field are quasinormal modes (QNMs), which are complex frequencies characterizing the oscillatory behavior of a black hole as it returns to a stable state after being disturbed. These modes are unique signatures that depend on the black hole’s intrinsic properties, such as its mass, charge, and spin, as well as the nature of the perturbation—whether it is scalar, electromagnetic, or gravitational. The independence of QNMs from initial conditions makes them particularly valuable in probing the characteristics of black holes, offering a direct link between theoretical predictions and observable phenomena Konoplya:2011qq ; Berti:2009kk ; Berti:2005ys ; Ferrari:2007dd ; Daghigh:2008jz ; Daghigh:2020jyk ; Daghigh:2020mog ; Daghigh:2006xg ; Daghigh:2022uws ; Fernando:2016ftj ; Fernando:2015kaa ; Fernando:2012yw ; Fernando:2008hb ; Chabab:2016cem ; Chabab:2017knz ; Lepe:2004kv ; Gonzalez:2017shu ; Khodadi:2018scn ; Gogoi:2022wyv ; Gogoi:2023kjt ; Gogoi:2021cbp ; Gogoi:2021dkr ; Oliveira:2021abg ; Kuang:2017cgt ; Kac:1966xd ; Andersson:1992scr ; Andersson:1994rm ; Andersson:1995vi ; Andersson:1996xw ; Maggio:2019zyv ; Konoplya:2022zym ; Konoplya:2018yrp ; Cardoso:2016rao ; Cardoso:2016oxy ; Visinelli:2017bny ; Vagnozzi:2020gtf ; Casalino:2018wnc ; Casalino:2018tcd ; Cardoso:2017cqb ; Vagnozzi:2023lwo ; Benetti:2021uea ; Yang:2019vni ; Pantig:2022gih ; Lambiase:2023hng ; Yang:2022ifo ; Okyay2022 ; Daghigh:2011ty ; Zhidenko:2003wq ; Zhidenko:2005mv ; Gonzalez:2022ote ; Rincon:2018ktz .
@@
The study of QNMs is especially critical in the context of the ringdown phase of binary black hole mergers. During this phase, the newly formed black hole emits gravitational waves as it stabilizes, with the QNMs encoding key information about the black hole’s parameters. As gravitational wave astronomy continues to grow, understanding QNMs has become increasingly important for interpreting the data from detectors like LIGO and Virgo LIGOScientific:2016aoc ; LIGOScientific:2016sjg .
@@
Our work is organized as follows: after the introduction we briefly review the CSBHs in Section II. In Section III., we study time-domain profile of the scalar field and then find the corresponding quasi-normal modes. In Section IV. and V., we investigate massless scalar and massless vector perturbations around the black hole. In Section VI. we study greybody factor using the WKB approach and also rigous bounds. Finally, we summarize and conclude our paper in Section VII.
@@ II Brief Review of Charged Symmergent Black Holes
Symmergent gravity is a specific form of emergent gravity that incorporates both $R$ and $R^{2}$ curvature terms along with a non-zero cosmological constant. As a specialized case within the broader class of $f(R)$ gravity theories, it was first introduced in dad2019 ; dad2016 , with further refinements and enhancements presented in dad2021 . This model has been applied across various contexts irfanaliben ; Symmergent-bh2 ; Pulice:2023dqw , while recent field-theoretic and string-theoretic developments have been detailed in dad2024 . In fact, symmergent gravity is defined by the action
@@
which is set by the mass-squared matrix ${\mathcal{M}}^{2}$ of the entirety of matter fields such that its supertrace ${\rm str}[{\mathcal{M}}^{2}]=\sum_{i}(-1)^{2s_{i}+1}(2s_{i}+1){\rm tr}[{\mathcal{M}}^{2}]_{s_{i}}$ runs over all fields $\psi_{i}$ with spin $s_{i}$. At the one-loop level, the quadratic curvature coefficient in Eq. (1) is given by the number difference between the total bosonic degrees of freedom $n_{B}$ and the fermionic degrees of freedom $n_{F}$ in the underlying QFT dad2023 , namely
@@
In the context of general $n_{\text{B}}$ and $n_{\text{F}}$, the Symmergent gravity action (1) may be expressed in the $f(R)$ gravity form as shown below Pulice:2023dqw :
@@
With the vacuum energy given by Eq. (11), the metric potential $h(r)$ takes the form Pulice:2023dqw :
@@
It results in the standard Reissner-Nordström-AdS/dS black hole when $\hat{Q}^{2}=\frac{Q^{2}}{2\alpha}$ and $\hat{\Lambda}=\frac{(1-\alpha)}{8\pi Gc_{\rm O}}$. The metric (10) with the potential (12) is called as CSBH Pulice:2023dqw .
@@
The predictions of the Einstein field equations have been tested extensively in a broad range of observationally accessible scenarios W:rev:14 , including but not limited to inflationary evolution inf1 , the formation of neutron stars and black holes irfanaliben ; Symmergent-bh2 ; reggiealiben ; bh4 ; bh5 ; cqgirfan ; Pulice:2023dqw ; Pulice:2024wjg , and the emission of gravitational waves gw1 ; gw2 . Given the increase in the number and variety of astrophysical observations of ultracompact objects (inlcuding recent EHT EventHorizonTelescope:2022xnr and JWST uncover observations), black holes manifest themselves as viable testbeds. This motivates our study of black hole spacetimes depending on the parameters $c_{\rm O}$ and $\alpha$.
III Scalar field propagation
How does a generic massless scalar field propagate within the given spacetime background? This leads us to the task of solving the massless Klein-Gordon equation in a fixed, curved background spacetime.
$\Box\Psi=0$ (13)
Given that the background spacetime is spherically symmetric and static, we can express $\Psi$ as a decomposition in terms of spherical harmonics,
$\Psi(t,r,\theta,\phi)=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}\frac{\psi_{lm}(r)}{r}e^{-i\omega t}Y_{lm}(\theta,\phi)\,.$ (14)
Substituting this ansatz into the Klein-Gordon equation yields the differential equation that governs the behavior of each individual mode
$h^{2}\frac{d^{2}\psi_{lm}}{dr^{2}}+hh^{\prime}\frac{d\psi_{lm}}{dr}+\omega^{2}\psi_{lm}=\frac{rhh^{\prime}+h(2l(l+1)+rh^{\prime})}{2r^{2}}\psi_{lm}\,.$ (15)
The equation can be further simplified by introducing the tortoise coordinate $r_{*}$, defined as:
$\frac{dr_{*}}{dr}$ $=h^{-1}$ (16)
Using the above definition of the tortoise coordinate $r_{*}$ and the details provided, we arrive at the following equation for $\psi_{lm}(r_{*})$ Berti:2009kk :
$\frac{d^{2}\psi_{lm}}{dr_{*}^{2}}+[\omega^{2}-V_{\text{eff}}(r_{*})]\psi_{lm}=0$ (17)
where $V_{\text{eff}}(r_{*})$, is the effective potential, depending on the specific form of $h(r)$, and the angular quantum numbers $l$ and $m$.
The effective potential $V_{\text{eff}}(r_{*})$, as a function of the tortoise coordinate $r_{*}$, is symmetric about $r_{*}=0$ (where $r$ corresponds to the horizon) and approaches to zero at large distances.
In the vicinity of the detector, the background spacetime can be approximated as flat. Here, we consider perturbations around a flat Minkowski background, along with a constant background scalar field. This allows us to treat the scalar field perturbations as small deviations from the flat spacetime, simplifying the analysis in this region
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu},\;\;\;\;\Phi=\Phi_{0}(1+\epsilon)$ (18)
where $\Phi_{0}$ is a constant. The field equations become
$\Box(-h_{\mu\nu}+\eta_{\mu\nu}\epsilon)=\frac{2{\bar{\kappa}}}{l\Phi_{0}}T_{\mu\nu}$ (19)
$\Box\epsilon=\frac{\bar{\kappa}}{l{\Phi}_{1}}T$ (20)
where ${\Phi}_{1}=\frac{3\Phi_{0}}{1+\Phi_{0}}$. Here the choice of gauge is
$\partial_{\nu}\bar{h}^{\mu\nu}=\partial^{\mu}\epsilon$ (21)
where ${\bar{h}}^{\mu\nu}$ is the trace-reversed metric perturbation. In vacuum, we have $\Box h_{\mu\nu}=0=\Box\epsilon$ and in the transverse traceless gauge, we get
$h_{\mu\nu}=\begin{bmatrix}0&0&0&0\\ 0&h_{+}-\epsilon_{0}&h_{\times}&0\\ 0&h_{\times}&-h_{+}-\epsilon_{0}&0\\ 0&0&0&0\end{bmatrix}e^{i\omega(t-z)}$ (22)
for a plane wave propagating in the $z$-direction. The scalar field is $\Phi=\Phi_{0}(1+\epsilon)$ where
$\epsilon=\epsilon_{0}e^{i\omega(t-z)}\,.$ (23)
The presence of the scalar field introduces an additional polarization in the gravitational wave, commonly referred to as the breathing mode Aneesh:2018hlp ; breathing .
Typically, if we consider a massive scalar field, a longitudinal mode would also emerge. However, in this work, we focus solely on a massless scalar field.
In a curved background, the equation governing the scalar field perturbation is given by:
$\Box(\delta\Phi)=g^{\mu\nu}\delta\Gamma^{\alpha}_{\mu\nu}\partial_{\alpha}\Phi+h^{\mu\nu}A_{\mu\nu}-B\,\delta\Phi-C^{\alpha}\partial_{\alpha}(\delta\Phi)$ (24)
where ()($\Box$)( □ ) represents the d’Alembertian operator in the curved spacetime and we have defined,
$A_{\mu\nu}$ $=$ $\frac{\Omega^{\prime}}{2\Omega+3}\Phi_{;\mu}\Phi_{;\nu}+\Phi_{;\mu;\nu}\,,$ (25)
$B$ $=$ $\frac{2\bar{\kappa}}{l}\frac{\Omega^{\prime}T^{b}}{(2\Omega+3)^{2}}+\frac{\Phi^{;\alpha}\Phi_{;\alpha}}{2\Omega+3}\left(\Omega^{\prime\prime}-\frac{2(\Omega^{\prime})^{2}}{2\Omega+3}\right)\,,$ (26)
$C^{\alpha}$ $=$ $\frac{2\Omega^{\prime}}{2\Omega+3}\Phi^{;\alpha}\,.$ (27)
Here, prime denotes differentiation with respect to $\Phi$, and we assume there is no fluctuation in $T^{b}$, the trace of the matter stress-energy tensor (i.e., $\delta T^{b}=0$), similarly in Zimmerman:2015hua . By performing an infinitesimal gauge transformation $h_{\mu\nu}\rightarrow h_{\mu\nu}+\xi_{\mu;\nu}+\xi_{\nu;\mu}$ and $\delta\Phi\rightarrow\delta\Phi+\xi^{\mu}\partial_{\mu}\Phi$, generated by $x^{\mu}\rightarrow x^{\mu}-\xi^{\mu}$, we obtain the following gauge condition:
$\left\{\Box\xi^{\alpha}-\xi^{\mu}R^{\alpha}_{\;\;\mu}-B\,\xi^{\alpha}\right\}\Phi_{;\alpha}+\xi^{\mu;\nu}\left(2A_{\mu\nu}-\Phi_{;\mu}C_{\nu}\right)-C^{\alpha}\Phi_{;\mu;\alpha}\xi^{\mu}=C^{\alpha}\partial_{\alpha}\delta\Phi-h^{\mu\nu}A_{\mu\nu}+B\,\delta\Phi-(h^{\alpha\nu}_{\,\,\,\,;\nu}-\frac{1}{2}h^{;\alpha})\Phi_{;\alpha}$ (28)
where $\xi^{\alpha}$ is the gauge function. With this, the equation for the scalar field perturbations simplifies to:
$\Box(\delta\Phi)=0$ (29)
Since the scalar perturbation satisfies the Klein-Gordon equation in a fixed background metric, the QNMs calculated correspond to the breathing mode mentioned earlier.
The metric and scalar field fluctuations generate the gravitational wave that is detected by instruments located in the asymptotic region, where the background spacetime can be approximated as flat. The scalar field fluctuation introduces a breathing mode polarization in the gravitational waves. The strain detected by a single detector is a combination of the projections of the gravitational waveforms associated with the different polarization states incident on it Isi:2017fbj . In general, it is challenging to isolate even the two polarization components predicted by the general relativity from a single detector’s observations Pai:2000zt , let alone the five non-degenerate polarization states allowed in metric theories of gravity Will:2018bme . However, with at least five linearly independent detectors, it becomes possible to resolve all five polarizations from transient signals Will:2018bme ; Sanjeev .
That solution, however, is beyond the scope of this work. Instead, we focus solely on the breathing mode and use the calculated QNMs to estimate errors on the metric parameters.
We now proceed to compute the time-domain profiles and the QNMs, which will serve as inputs for parameter estimation using gravitational wave data.
IV Perturbations and Quasinormal modes

### v2

@@ Abstract
In this paper, we investigate quasinormal modes (QNMs) and greybody factors within the framework of Symmergent gravity, an emergent gravity model with an $R+R^{2}$ curvature sector. Building on our previous work on static spherically-symmetric solutions [Class.Quant.Grav. 40 (2023) 19, 195003], we explore the effects of the key parameters, including the quadratic curvature coupling parameter $c_{\rm O}$ and the vacuum energy parameter $\alpha$. For both scalar and electromagnetic perturbations, an increase in $\alpha$ leads to a nearly linear rise in both oscillation frequencies and damping rates. The other parameter $c_{\rm O}$ affects the QNMs spectrum nonlinearly. Additionally, the charge $Q$ of the black hole introduces nonlinear behavior, where higher charges amplify the black hole’s electromagnetic field, resulting in increased oscillation frequencies and faster stabilization. These findings enhance our understanding of charged black hole stability and gravitational wave astrophysics. Further, the analysis of greybody factors reveals that increasing $\alpha$, $c_{\rm O}$, and $Q$ reduces the absorption of radiation, with electromagnetic perturbations reaching maximum absorption at slightly lower frequencies compared to scalar perturbations.
@@ I Introduction
The Standard Model (SM) describes the strong, weak, and electromagnetic interactions as a renormalizable quantum field theory (QFT) [1], but does not account for gravity. It is unique to the flat Minkowski spacetime because of difficulties in quantizing the curved metric [2] and in transferring QFTs into curved spacetime [3, 4]. In contrast to full QFTs like the SM, flat spacetime effective QFTs are expected to exhibit a closer connection to curved spacetime. This is because they are nearly classical field theories obtained by integrating out high-frequency quantum fluctuations [1, 5], in the framework of both Wilsonian and one-particle-irreducible effective actions [7, 8].
@@
The SM is endowed with a physical UV cutoff (rather than a regulator) when viewed as an effective QFT [5]. This UV cutoff denoted by $\Lambda_{\wp}$, represents the scale at which a UV completion becomes relevant. As revealed by the null LHC searches [6], $\Lambda_{\wp}$ could be at any scale above a TeV. In general, in a Lorentz conserving regime, the loop momentum $\ell^{\mu}$ lies in the interval $-\Lambda_{\wp}^{2}\leq\eta_{\mu\nu}\ell^{\mu}\ell^{\nu}\leq\Lambda_{\wp}^{2}$, and it explicitly breaks the Poincaré (translation) symmetry. In the presence of the UV cutoff $\Lambda_{\wp}^{2}$, scalar mass-squareds obtain corrections proportional to $\Lambda_{\wp}^{2}$, and $\Lambda_{\wp}^{2}$ and $\Lambda_{\wp}^{4}$ terms appear in the loop-corrected vacuum energy [5, 9]. Beyond these UV sensitivities [5], gauge bosons also acquire mass-squared terms proportional to $\Lambda_{\wp}^{2}$, and which leads to a breaking of gauge symmetries [10, 11]. In order to alleviate these unnatural UV sensitivities, the most sensible line of action would be to restore the gauge symmetries by means of the Higgs mechanism [12, 13, 14]. A hindrance to this proposal is the stark difference between the intermediate vector boson mass (Poincaré-conserving) [15] and the loop-induced gauge boson mass (Poincaré-breaking) [16]. Indeed, in the former, the vector boson mass is promoted to a scalar field, which leads to the usual Higgs mechanism [12, 13, 14]. In the latter, however, it is necessary to first find a Poincaré-breaking Higgs field. In this regard, it turns out that the affine curvature, which is independent of the metric tensor and its connection, is the sought-after Higgs field in that it promotes the UV cutoff $\Lambda_{\wp}$ [17, 18] and condenses to the usual metrical curvature in the minimum of the metric-affine action [19, 20]. The defusion of the unnatural UV sensitivity, the restoration of gauge symmetries, the emergence of gravity, and the appearance of new particles without the necessity to couple directly to SM particles are the salient outcomes of this condensation. This whole mechanism is referred to as gauge symmetry-restoring emergent gravity or symmergent gravity in brief. Symmergent gravity has been constructed and applied through various stages [16, 22, 23, 21], starting with the nascent idea of gauge symmetry restoration by means of curvature [24]. Some applications of Symmergent gravity can be found in [30, 31, 26, 27, 25, 31, 32, 34, 33, 29].
@@
Black holes are not isolated cosmic bodies; rather, they are dynamic entities that actively interact with their surroundings, profoundly influencing the fabric of spacetime. When a black hole is perturbed, spacetime itself ripples, creating waves that eventually die down as the system settles back to equilibrium. Understanding how these systems respond to small disturbances is a fundamental aspect of physics, providing deep insights into the nature of spacetime and gravity. The study of black hole perturbations began with pioneering work by [42], and was significantly advanced by others [43, 44, 45, 46, 47, 48], leading to Chandrasekhar’s seminal contributions [49, 50]. At the heart of this field are quasinormal modes (QNMs), which are complex frequencies characterizing the oscillatory behavior of a black hole as it returns to a stable state after being disturbed. These modes are unique signatures that depend on the black hole’s intrinsic properties, such as its mass, charge, and spin, as well as the nature of the perturbation—whether it is scalar, electromagnetic, or gravitational. The independence of QNMs from initial conditions makes them particularly valuable in probing the characteristics of black holes, offering a direct link between theoretical predictions and observable phenomena [51, 52, 58, 59, 53, 54, 55, 56, 57, 64, 65, 66, 67, 60, 61, 62, 63, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101].
@@
The study of QNMs is especially critical in the context of the ringdown phase of binary black hole mergers. During this phase, the newly formed black hole emits gravitational waves as it stabilizes, with the QNMs encoding key information about the black hole’s parameters. As gravitational wave astronomy continues to grow, understanding QNMs has become increasingly important for interpreting the data from detectors like LIGO and Virgo [102, 103].
@@
Our work is organized as follows: after the introduction, we briefly review the CSBHs in Section II. In Sections III and IV, we investigate massless scalar and massless vector perturbations around the black hole. In Section V, we study greybody factor using the WKB approach and rigorous bounds. Finally, we summarize and conclude our paper in Section VI.
@@ II Brief Review of Charged Symmergent Black Holes
Symmergent gravity is a specific form of emergent gravity that incorporates both $R$ and $R^{2}$ curvature terms along with a non-zero cosmological constant. As a specialized case within the broader class of $f(R)$ gravity theories, it was first introduced in [23, 24], with further refinements and enhancements presented in [22]. This model has been applied across various contexts [30, 31, 27], while recent field-theoretic and string-theoretic developments have been detailed in [21]. In fact, symmergent gravity is defined by the action
@@
which is set by the mass-squared matrix ${\mathcal{M}}^{2}$ of the entirety of matter fields such that its supertrace ${\rm str}[{\mathcal{M}}^{2}]=\sum_{i}(-1)^{2s_{i}+1}(2s_{i}+1){\rm tr}[{\mathcal{M}}^{2}]_{s_{i}}$ runs over all fields $\psi_{i}$ with spin $s_{i}$. At the one-loop level, the quadratic curvature coefficient in Eq. (1) is given by the number difference between the total bosonic degrees of freedom $n_{B}$ and the fermionic degrees of freedom $n_{F}$ in the underlying QFT [16], namely
@@
In the context of general $n_{\text{B}}$ and $n_{\text{F}}$, the Symmergent gravity action (1) may be expressed in the $f(R)$ gravity form as shown below [27]:
@@
With the vacuum energy given by Eq. (11), the metric potential $h(r)$ takes the form [27]:
@@
It results in the standard Reissner-Nordström-AdS/dS black hole when $\hat{Q}^{2}=\frac{Q^{2}}{2\alpha}$ and $\hat{\Lambda}=\frac{(1-\alpha)}{8\pi Gc_{\rm O}}$. The metric (10) with the potential (12) is called as CSBH [27].
@@
The predictions of the Einstein field equations have been tested extensively in a broad range of observationally accessible scenarios [28], including but not limited to inflationary evolution [29], the formation of neutron stars and black holes [30, 31, 32, 33, 35, 34, 27, 26], and the emission of gravitational waves [36, 37]. Given the increase in the number and variety of astrophysical observations of ultracompact objects (inlcuding recent EHT [38] and JWST [39] observations), black holes manifest themselves as viable testbeds. This motivates our study of black hole spacetimes depending on the parameters $c_{\rm O}$ and $\alpha$.
III Perturbations and Quasinormal modes